OpenStudy (anonymous):

2x + y = 1, x + 2y = 1

OpenStudy (anonymous):

x + 2y = 1 x = 1 - 2y Plug it in instead of x in the first equation and solve for x.

OpenStudy (anonymous):

so y=1?

OpenStudy (anonymous):

Do you need to find x and y? If so, here are the steps... x+2y=1 x=-2y+1 (Plug that into x+2y=1) -2y+1+2y=1 Then figure it out from there.